Five grams of sugar are added to a 45-gram serving of a breakfast cereal that is 20% sugar. What is the percent concentration of sugar in the resulting mixture?
? %

Answer by stanbon(75887)

Five grams of sugar are added to a 45-gram serving of a breakfast cereal that is 20% sugar. What is the percent concentration of sugar in the resulting mixture?
--------------
Equation:
sugar + sugar = sugar
0.20*45 + 5 = x(45+5)
---
Multiply thru by 100 to get:
20*45 + 500 = 100x(50)
1400 = 5000x
x = 0.28
----
Ans: 28%
=============
Cheers,
Stan H.